# Load Forecasting for Households and Energy Communities: Are Deep Learning Models Worth the Effort?

### v1

@@ Abstract
keywords:
Load Forecasting , Deep Learning , Transformer , xLSTM , Transfer Learning , Energy Communities
journal: Energy and AI
[…]
@@ 1 Introduction
For the management of energy assets, model predictive control (MPC) methods are used with great success 1, 2. In practice, MPC relies on accurate forecasts for exogenous variables such as load, production, or emissions. Thus, forecasting short-term load is a crucial part for the optimized operation of flexible energy assets 3.
@@
Increasingly popular entities, concerned with the control of flexible assets, are Energy Communities (ECs). ECs are groups of individuals or organizations that produce, manage, and consume energy locally, through the use of public grid infrastructure, often with the goal of increasing sustainability and reducing energy costs 4. Efficient control of their energy assets through, for example, MPC can increase their financial benefit5.
@@
Even though this works’ focus is on ECs, aggregated household load forecasting can be applied in other applications as well. While small ECs are similar to the load behaviors of individual households, larger ECs are similar to the load profiles of secondary substations. Numerous models have already been proposed for short-term load forecasting, covering time horizons ranging from one hour up to one week 3. In recent years, neural networks using Long Short-Term Memory (LSTM) have been widely regarded as popular models for short-term load forecasting of households 3, 6, 7. However, more recently, many studies have replaced LSTMs with Transformer models due to their ability to handle long-range dependencies and efficiently parallelize computations. For example, Semmelmann et al. 8 compared the forecasting errors of LSTMs and Transformers for an EC consisting of 21 households and concluded that Transformer models outperformed LSTMs. Similarly, Ran et al. 9 achieved the same result when evaluating a single highly aggregated load profile representing the entire city of New York. Likewise, Zhao et al. 10 found that Transformers significantly outperformed LSTMs, using data from a single aggregated load profile representing New South Wales, Australia. L’Heureux et al. 11 showed that a Transformer-based model outperformed a state-of-the-art recurrent model when tested on 20 highly aggregated load profiles from an energy utility company. Fang et al. 12 also reported similar findings, where a Transformer-based approach surpassed an LSTM-based model using a single high-aggregation load profile from one zone of Tetouan city.
@@
In May 2024, Beck et al. 13 introduced the xLSTM model. The authors claimed that this model performs favorably when compared to state-of-the-art Transformers, both in performance and scaling. They demonstrated that xLSTM exhibits performance comparable to Transformers, with notable advantages such as linear rather than quadratic growth in memory consumption as the model input size increases. This improvement addresses scalability challenges, making xLSTM a more efficient option to handle larger inputs. A first work by Kraus et al. 14 demonstrates the effectiveness of xLSTM for long-term forecasting across various publicly available time series, with forecast horizons ranging from 96 to 720 hours, in contrast to the short-term forecasting problem addressed in this study.
@@
One practical problem of such deep learning models in ECs is that they often lack enough historical data 15. This potentially leads to poor system performance. Transfer learning is a promising approach to overcome this problem. This approach is based on the idea that knowledge learned during one task can be reused to boost the performance on a related task 16. In Moosbrugger et al.17, the potential of leveraging transfer learning was demonstrated using open-access synthetic (also called standard) load profiles to improve load forecast error for renewable ECs with limited historical data. The study utilized a bidirectional LSTM model pre-trained on open-access synthetic profiles, achieving a notable reduction in mean squared error (MSE) from 0.34 to 0.13 in a test EC involving 74 households. Furthermore, this approach significantly improved training stability.
@@
Compared to the load forecasting studies mentioned, this research is broader in scope. Profiles, aggregation levels, training data durations, seasons, and model sizes are varied, providing a more comprehensive evaluation. This approach aligns with the recommendation of Pinheiro et al.18, who noted that many existing studies demonstrate the superiority of a proposed technique on narrowly defined datasets, often hiding their weaknesses. By introducing diverse setups, this study seeks to reflect realistic variations, providing insights that are both more generalizable and practically applicable. In addition, a case study is presented that quantifies the financial implications of different forecast errors. In the case study, MPC is applied to a simulated EC with different forecasting algorithms.
@@ 2 Methods
To assess the performance of all models for the various given use cases, a software framework is designed and implemented, detailed in this Section. Specifically, three deep learning models - xLSTM, LSTM, and Transformer - are compared against three basic models: Persistence Prediction, k-Nearest Neighbors (KNN), and forecasting with synthetic load. Each model performs day-ahead short-term load forecasting, generating 24 hourly load predictions spanning from midnight to midnight. The xLSTM, LSTM, Transformer, and KNN models utilize standard input features for short-term load forecasting, including datetime information, weather conditions, and historical load data, as utilized by Pinheiro et al. 18.
The models are trained and tested on various configurations of ECs, aggregation sizes, training dataset sizes, seasons, and model parameter counts. A parameter grid is constructed, and a baseline setup is defined. This baseline allows to systematically analyze how forecast errors vary when specific parameters are individually adjusted. This enables to manage the complexity of the analysis while maintaining a focus on the most critical factors influencing model performance. More documentation, the full implementation, and the datasets used can be found in the corresponding GitHub repository 19.
@@ 2.1.1 Aggregation Size
To investigate different levels of aggregation, five different community sizes have been selected for comparison. The Austrian Coordination Office for Energy Communities 20 categorizes most ECs in Austria into 2–10, 10–50, or 50–100 households. Based on this classification, the boundary cases of 1, 2, 10, 50, and 100 households are selected for analysis. In the parameter grid, 50 households are chosen as the baseline aggregation size, representing a central point between 1 and 100. This baseline serves as a reference point for systematically evaluating the forecast error.
@@ 2.2 Dataset
The dataset used in this study is derived from the smart meter energy consumption data provided by the UK Power Networks uk_power_networks_2015, which includes records from 5,567 households in London. To construct virtual ECs, subsets of 2, 10, 50, or 100 households were randomly selected from the refined dataset. This selection process was repeated 20 times for each category of community size, without replacement, resulting in $20\cdot 4=80$ unique virtual ECs, along with 20 individual household profiles. All profiles span the period from July 2012 to February 2014, providing a robust foundation for the analysis.
@@ 2.3.1 Basic Models (Non Deep Learning)
Persistent Prediction Model: This benchmark model predicts the next day to be exactly the same as it was 7 days ago. Gross et al. 21 call this also the naive seasonal model.
KNN21: This model utilizes the same input features as the deep learning models, described in 2.3.3. In the implementation of this study, k was set to 1, meaning the model predicts the next day based on the day with the most similar features according to Euclidean distance.
@@ 2.3.2 Deep Learning Models
Two very popular deep learning methods for load forecasting are selected from the literature review 3, 6, 7, 8, 9, 10, 11, 12 for this comparative study : LSTMs and Transformers. Additionally, a new model, xLSTM13, is chosen.
The general model architecture used in this study is illustrated in Fig. 1. Similar to many established LSTM-based applications (see, e.g., Kong et al.6), the LSTM layer is followed by multiple dense (i.e., feedforward neural network) layers. We use the same approach for xLSTMs. For the Transformer model, an encoder-only architecture is adopted, as demonstrated by Nie et al. 22. To ensure fair comparability across the different sequence-to-sequence deep learning models, the dense layers at the output are designed with a fixed number of neurons.
@@
All models are trained with Adam 23 optimizer over 100 epochs, with dynamic learning rate from 0.01 to 0.0005 and using batch sizes of 256.
@@ 2.3.3 Deep Learning Input Features
The current datetime, past weather data, and past load values are used as input features, which are, for instance, also used by Pinheiro et al.18. These hourly model input features are detailed as follows:
  • 1.
@@
  • 5.
    Inputs 14-19: Weather data. All available Meteostat 24 weather features for London from 2011 to 2014 are utilized. These features include temperature, dew point, wind direction, wind speed, air pressure, and relative humidity from the last 24 hours before the forecast.
@@ 2.3.4 Model Size
When tuning hyperparameters of these deep learning models, the aim is to ensure that the model sizes (measured by parameter count) are roughly comparable across architectures. Similar to the test set selection process described in Section 2.1.3, preliminary analyses are conducted to identify effective model sizes. The results of these experiments (presented in ion 3.3) indicate that a target model size of 5k parameters serves as a good baseline. This size achieved robust performance across all models and lies at the midpoint of the parameter range, balancing model complexity and forecast error.
@@ 2.3.5 Transfer Learning
Building on the authors’ previous work 17, transfer learning is applied using synthetic load profiles. During pretraining, models are exposed to synthetic load profiles spanning the entire period from July 2012 to February 2014. The learned parameters from this phase are saved and later loaded before fine-tuning the model with real EC profiles. The results were generated with and without transfer learning. The baseline for the parameter grid includes transfer learning.
@@ 2.4 Implementation and Evaluation
The framework is evaluated on an Ubuntu 22.04 system. To ensure reproducibility, a conda environment file is supplied, detailing the primary dependencies, including Python 3.11 25 and PyTorch 2.2 26.
@@ 4 Case Study
This section presents a simulation case study to quantify the financial implications of forecasting errors within an EC. The objective is to reduce daily energy costs via intraday load shift. The EC consists of 50 households and a community battery energy storage system (BESS) assumed to be located in Austria. It has a real-time price (RTP) tariff based on the European Power Exchange (EPEX) 27. In this case study, the Energy Exchange Austria (EXAA) 28 is used in combination with network usage fees and taxes according to current Austrian standards. A 4-hour (C-Rate = 0.25) BESS with a capacity of 1 MWh with a round-trip efficiency of 85% 29 is installed. The BESS is used for load shifting, to minimize the EC electricity costs. The simulation is executed for the test period from 1st October 2013 to 31st December 2013 and the simulation is discretized to hourly time steps $\Delta t$.
@@
$E,P_{\text{ch}},_{\text{dis}},\hat{P}_{\text{grid},p}\in\mathbb{R}$ (14)
@@
EEq. (3) defines the objective function minimizing the daily electrical energy costs. Eq. (4) defines a set of indices for all time periods and Eq. (5) defines a set of indices for all time points. The power balance of the EC is defined in Eq. (6). Eqs. (7-11) describe the BESS model, where, $\eta_{\text{ch}}$ and $\eta_{\text{dis}}$ are the charging and discharging efficiencies, respectively and $E_{\text{min}}$ and $E_{\text{max}}$ denote minimum and maximum BESS capacity, respectively. The energy balance of the BESS is described in Eq. (7). Eqs. (8-10) prevent simultaneous charging and discharging via the binary variables $b_{\text{ch}}$ and $b_{\text{dis}}$. The initial and the end conditions of the BESS are defined by Eqs. (12,13).t the optimization uses the forecasteded load $P_{\text{load}}$ without a feedback loop, as real-time measurements over 50 households can be considered very difficult in practice. Consequently, the grid is used to compensate forecast errors. The real, resulting grid power $P_{\text{grid}}$ can therefore be described as such:
@@
Besides perfect prediction, which assumes perfect knowledge of the signal, the best results are achieved using the Transformer leading to savings of 1303.85€, which is equal to a cost reduction of 10.20%. LSTM and xLSTM perform nearly as well as Transformer with a cost reduction of 9.94% for LSTM and 9.95% for xLSTM. This shows that all three deep learning algorithms perform well and achieve savings of up to 161.02€ompared with persistent prediction over the investigated three months. For one year, these savings would scale to approximately 600€. In this case study, all prediction methods perform relatively well, because for RTP-driven load shifting, the price signal is the most important factor which is published one day in advance by EXAA. In contrast, other objectives, such as increasing self-sufficiency or reducing peak load are more prediction-sensitive, emphasizing the significance of high-quality predictions30. This study demonstrates the suitability of deep learning for predicting electrical loads in an EC, with significant potential for load shifting to reduce operational costs.
@@ 5 Conclusions
One of the main findings of this study is that ersistence prediction performs best when an EC has only six or fewer months of training data. However, as more data become available, models achieve lower forecast errors.
@@
A case study quantifies the financial implication of the improvement between deep learning models and persistence prediction. The difference in forecast error of approximately 1.1%pt to 600€ annual savings for an EC of 50 households with a battery storage.
@@ Appendix A Model Configuration
Layers = 1 Heads = 2 dim_feedforward = 10 Features = 16

### v2

@@ 1 Introduction
For the management of energy assets, model predictive control (MPC) methods are used with great success [1, 2]. In practice, MPC relies on accurate forecasts for exogenous variables such as load, production, or emissions. Thus, forecasting short-term load is a crucial part for the optimized operation of flexible energy assets [3].
@@
Increasingly popular entities, concerned with the control of flexible assets, are Energy Communities (ECs). ECs are groups of individuals or organizations that produce, manage, and consume energy locally, through the use of public grid infrastructure, often with the goal of increasing sustainability and reducing energy costs [4]. Efficient control of their energy assets through, for example, MPC can increase their financial benefit[5].
Even though this works’ focus is on ECs, aggregated household load forecasting can be applied in other applications as well. While small ECs are similar to the load behaviors of individual households, larger ECs are similar to the load profiles of secondary substations.
@@
Numerous models have already been proposed for short-term load forecasting, covering time horizons ranging from one hour up to one week [3]. In recent years, neural networks using Long Short-Term Memory (LSTM) have been widely regarded as popular models for short-term load forecasting of households [3, 6, 7]. However, more recently, many studies have replaced LSTMs with Transformer models due to their ability to handle long-range dependencies and efficiently parallelize computations. For example, Semmelmann et al. [8] compared the forecasting errors of LSTMs and Transformers for an EC consisting of 21 households and concluded that Transformer models outperformed LSTMs. Similarly, Ran et al. [9] achieved the same result when evaluating a single highly aggregated load profile representing the entire city of New York. Likewise, Zhao et al. [10] found that Transformers significantly outperformed LSTMs, using data from a single aggregated load profile representing New South Wales, Australia. L’Heureux et al. [11] showed that a Transformer-based model outperformed a state-of-the-art recurrent model when tested on 20 highly aggregated load profiles from an energy utility company. Fang et al. [12] also reported similar findings, where a Transformer-based approach surpassed an LSTM-based model using a single high-aggregation load profile from one zone of Tetouan city.
@@
In May 2024, Beck et al. [13] introduced the xLSTM model. The authors claimed that this model performs favorably when compared to state-of-the-art Transformers, both in performance and scaling. They demonstrated that xLSTM exhibits performance comparable to Transformers, with notable advantages such as linear rather than quadratic growth in memory consumption as the model input size increases. This improvement addresses scalability challenges, making xLSTM a more efficient option to handle larger inputs. A first work by Kraus et al. [14] demonstrates the effectiveness of xLSTM for long-term forecasting across various publicly available time series, with forecast horizons ranging from 96 to 720 hours, in contrast to the short-term forecasting problem addressed in this study.
@@
One practical problem of such deep learning models in ECs is that they often lack enough historical data [15]. This potentially leads to poor system performance. Transfer learning is a promising approach to overcome this problem. This approach is based on the idea that knowledge learned during one task can be reused to boost the performance on a related task [16]. In Moosbrugger et al.[17], the potential of leveraging transfer learning was demonstrated using open-access synthetic (also called standard) load profiles to improve load forecast error for renewable ECs with limited historical data. The study utilized a bidirectional LSTM model pre-trained on open-access synthetic profiles, achieving a notable reduction in mean squared error (MSE) from 0.34 to 0.13 in a test EC involving 74 households. Furthermore, this approach significantly improved training stability.
@@
Compared to the load forecasting studies mentioned, this research is broader in scope. Profiles, aggregation levels, training data durations, seasons, and model sizes are varied, providing a more comprehensive evaluation. This approach aligns with the recommendation of Pinheiro et al.[18], who noted that many existing studies demonstrate the superiority of a proposed technique on narrowly defined datasets, often hiding their weaknesses. By introducing diverse setups, this study seeks to reflect realistic variations, providing insights that are both more generalizable and practically applicable. In addition, a case study is presented that quantifies the financial implications of different forecast errors. In the case study, MPC is applied to a simulated EC with different forecasting algorithms.
@@ 2 Methods
To assess the performance of all models for the various given use cases, a software framework is designed and implemented, detailed in this Section. Specifically, three deep learning models - xLSTM, LSTM, and Transformer - are compared against three basic models: Persistence Prediction, k-Nearest Neighbors (KNN), and forecasting with synthetic load. Each model performs day-ahead short-term load forecasting, generating 24 hourly load predictions spanning from midnight to midnight. The xLSTM, LSTM, Transformer, and KNN models utilize standard input features for short-term load forecasting, including datetime information, weather conditions, and historical load data, as utilized by Pinheiro et al. [18].
The models are trained and tested on various configurations of ECs, aggregation sizes, training dataset sizes, seasons, and model parameter counts. A parameter grid is constructed, and a baseline setup is defined. This baseline allows to systematically analyze how forecast errors vary when specific parameters are individually adjusted. This enables to manage the complexity of the analysis while maintaining a focus on the most critical factors influencing model performance. More documentation, the full implementation, and the datasets used can be found in the corresponding GitHub repository [19].
@@ 2.1.1 Aggregation Size
To investigate different levels of aggregation, five different community sizes have been selected for comparison. The Austrian Coordination Office for Energy Communities [20] categorizes most ECs in Austria into 2–10, 10–50, or 50–100 households. Based on this classification, the boundary cases of 1, 2, 10, 50, and 100 households are selected for analysis. In the parameter grid, 50 households are chosen as the baseline aggregation size, representing a central point between 1 and 100. This baseline serves as a reference point for systematically evaluating the forecast error.
@@ 2.2 Dataset
The dataset used in this study is derived from the smart meter energy consumption data provided by the UK Power Networks [21], which includes records from 5,567 households in London. To construct virtual ECs, subsets of 2, 10, 50, or 100 households were randomly selected from the refined dataset. This selection process was repeated 20 times for each category of community size, without replacement, resulting in $20\cdot 4=80$ unique virtual ECs, along with 20 individual household profiles. All profiles span the period from July 2012 to February 2014, providing a robust foundation for the analysis.
@@ 2.3.1 Basic Models (Non Deep Learning)
Persistent Prediction Model: This benchmark model predicts the next day to be exactly the same as it was 7 days ago. Gross et al. [22] call this also the naive seasonal model.
KNN[22]: This model utilizes the same input features as the deep learning models, described in 2.3.3. In the implementation of this study, k was set to 1, meaning the model predicts the next day based on the day with the most similar features according to Euclidean distance.
@@ 2.3.2 Deep Learning Models
Two very popular deep learning methods for load forecasting are selected from the literature review [3, 6, 7, 8, 9, 10, 11, 12] for this comparative study : LSTMs and Transformers. Additionally, a new model, xLSTM[13], is chosen.
The general model architecture used in this study is illustrated in Fig. 1. Similar to many established LSTM-based applications (see, e.g., Kong et al.[6]), the LSTM layer is followed by multiple dense (i.e., feedforward neural network) layers. We use the same approach for xLSTMs. For the Transformer model, an encoder-only architecture is adopted, as demonstrated by Nie et al. [23]. To ensure fair comparability across the different sequence-to-sequence deep learning models, the dense layers at the output are designed with a fixed number of neurons.
@@
All models are trained with Adam [24] optimizer over 100 epochs, with dynamic learning rate from 0.01 to 0.0005 and using batch sizes of 256.
@@ 2.3.3 Deep Learning Input Features
The current datetime, past weather data, and past load values are used as input features, which are, for instance, also used by Pinheiro et al.[18]. These hourly model input features are detailed as follows:
@@
  • Inputs 14-19: Weather data. All available Meteostat [25] weather features for London from 2011 to 2014 are utilized. These features include temperature, dew point, wind direction, wind speed, air pressure, and relative humidity from the last 24 hours before the forecast.
@@ 2.3.4 Model Size
When tuning hyperparameters of these deep learning models, the aim is to ensure that the model sizes (measured by parameter count) are roughly comparable across architectures. Similar to the test set selection process described in Section 2.1.3, preliminary analyses are conducted to identify effective model sizes. The results of these experiments (presented in Section 3.3) indicate that a target model size of 5k parameters serves as a good baseline. This size achieved robust performance across all models and lies at the midpoint of the parameter range, balancing model complexity and forecast error.
@@ 2.3.5 Transfer Learning
Building on the authors’ previous work [17], transfer learning is applied using synthetic load profiles. During pretraining, models are exposed to synthetic load profiles spanning the entire period from July 2012 to February 2014. The learned parameters from this phase are saved and later loaded before fine-tuning the model with real EC profiles. The results were generated with and without transfer learning. The baseline for the parameter grid includes transfer learning.
@@ 2.4 Implementation and Evaluation
The framework is evaluated on an Ubuntu 22.04 system. To ensure reproducibility, a conda environment file is supplied, detailing the primary dependencies, including Python 3.11 [26] and PyTorch 2.2 [27].
@@ 4 Case Study
This section presents a simulation case study to quantify the financial implications of forecasting errors within an EC. The objective is to reduce daily energy costs via intraday load shift. The EC consists of 50 households and a community battery energy storage system (BESS) assumed to be located in Austria. It has a real-time price (RTP) tariff based on the European Power Exchange (EPEX) [28]. In this case study, the Energy Exchange Austria (EXAA) [29] is used in combination with network usage fees and taxes according to current Austrian standards. A 4-hour (C-Rate = 0.25) BESS with a capacity of 1 MWh with a round-trip efficiency of 85% [30] is installed. The BESS is used for load shifting, to minimize the EC electricity costs. The simulation is executed for the test period from 1st October 2013 to 31st December 2013 and the simulation is discretized to hourly time steps $\Delta t$.
@@
$E,P_{\text{ch}},P_{\text{dis}},\hat{P}_{\text{grid},p}\in\mathbb{R}$ (14)
@@
Eq. (3) defines the objective function minimizing the daily electrical energy costs. Eq. (4) defines a set of indices for all time periods and Eq. (5) defines a set of indices for all time points. The power balance of the EC is defined in Eq. (6). Eqs. (7-11) describe the BESS model, where, $\eta_{\text{ch}}$ and $\eta_{\text{dis}}$ are the charging and discharging efficiencies, respectively and $E_{\text{min}}$ and $E_{\text{max}}$ denote minimum and maximum BESS capacity, respectively. The energy balance of the BESS is described in Eq. (7). Eqs. (8-10) prevent simultaneous charging and discharging via the binary variables $b_{\text{ch}}$ and $b_{\text{dis}}$. The initial and the end conditions of the BESS are defined by Eqs. (12,13). Note that the optimization uses the forecasted load $P_{\text{load}}$ without a feedback loop, as real-time measurements over 50 households can be considered very difficult in practice. Consequently, the grid is used to compensate forecast errors. The real, resulting grid power $P_{\text{grid}}$ can therefore be described as such:
@@
Besides perfect prediction, which assumes perfect knowledge of the signal, the best results are achieved using the Transformer leading to savings of 1303.85€, which is equal to a cost reduction of 10.20%. LSTM and xLSTM perform nearly as well as Transformer with a cost reduction of 9.94% for LSTM and 9.95% for xLSTM. This shows that all three deep learning algorithms perform well and achieve savings of up to 161.02€ compared with persistent prediction over the investigated three months. For one year, these savings would scale to approximately 600€. In this case study, all prediction methods perform relatively well, because for RTP-driven load shifting, the price signal is the most important factor which is published one day in advance by EXAA. In contrast, other objectives, such as increasing self-sufficiency or reducing peak load are more prediction-sensitive, emphasizing the significance of high-quality predictions[31]. This study demonstrates the suitability of deep learning for predicting electrical loads in an EC, with significant potential for load shifting to reduce operational costs.
@@ 5 Conclusions
One of the main findings of this study is that persistence prediction performs best when an EC has only six or fewer months of training data. However, as more data become available, deep learning models achieve lower forecast errors.
@@
A case study quantifies the financial implication of the improvement between deep learning models and persistence prediction. The difference in forecast error of approximately 1.1%pt can be translated to 600€ annual savings for an EC of 50 households with a battery storage.
@@ Appendix A Model Configuration
Layers = 1
Heads = 2
dim_feedforward = 10
Features = 16